The distance around a room is 24 M. The room is twice as long as it is wide what is the length and width

Answer by addingup(3677)

2L + 2W = 24
L = 2W
2(2W) + 2W = 24
4W + 2W = 24
6W = 24
W = 4
---------------------
Check:
W = 4
L = 2W = 2(4) = 8
2(4) + 2(8) = 24
8 + 16 = 24 correct